# Constraining the abundance of Galactic compact objects with continuous gravitational waves

### v1

@@ Abstract
Galactic spinning compact objects (COs) with non-zero ellipticity are expected to be sources of continuous gravitational waves (CGWs). Certain classes of hypothetical COs, such as neutron stars with quark cores (hybrid stars), and quark stars, are thought to be capable of sustaining large ellipticities from theoretical considerations. Such exotic COs (eCOs) with large ellipticities should produce CGWs detectable by the current LIGO-Virgo-Kagra GW detector network. Since no detections for CGWs, from searches in LIGO-Virgo data, have so far been reported, we place constraints on the abundance of highly elliptical eCOs in our Galaxy. We formulate a Bayesian framework to place upper limits on the number count $N_{tot}$ of highly deformed Galactic eCOs. We divide our constraints into two classes: an “agnostic” set of upper limits on $N_{tot}$ evaluated on a CGW frequency and ellipticity grid that depend only on the choice of spatial distribution of COs; and a model-dependent set that additionally assumes prior information on the distribution of frequencies. We find that COs with ellipticities $\epsilon\gtrsim 10^{-5}$ have abundance upper limits at $90\%$ confidence, of $N_{tot}^{90\%}\lesssim 100$, and those with $\epsilon\gtrsim 10^{-6}$ have $N_{tot}^{90\%}\lesssim 10^{4}$. We additionally place upper-limits on the ellipticity of Galactic COs informed by our choices of spatial distributions, given different abundances $N_{tot}$.
@@ 1 Introduction
Conversely, the reported non-detection of CGWs from all-sky searches in LV’s O1, O2, O3 data, has strong implications on the abundance of Galactic eCOs. In this letter, we formulate a Bayesian framework to constrain their number count. Specifically, from the CGW (frequency-dependent and sky-averaged) amplitude upper limits (Abbott et al., 2022d), and assuming that the eCOs are Poisson distributed, we construct a Bayesian posterior distribution on the number of COs ($N_{tot}$) from which we place upper limits on the number count at $90\%$ confidence.
@@
We additionally place spatial-distribution-informed upper limits on the ellipticity, assuming CO number counts of $10^{7,8,9}$ (Treves et al., 2000), consistent with the expected abundance of Galactic NSs. Our most conservative estimate yields $\epsilon\lesssim 1.5\times 10^{-7}$.
@@ 2.4 Bayesian framework
Figure 3: Histograms of samples drawn from the posterior on $N_{tot}$ (cf. Eq. 12), marginalized with respect to the ATNF-like frequency distribution. The progenitor and $3$ “Reed” distributions are considered, and the histograms are evaluated for four choices of ellipticities $\epsilon=10^{-7,-6,-5,-4}$. The upper limits for $\epsilon=10^{-5,-4}$ are less than $100$, suggesting that the abundance of such high-ellipticity objects in the Galaxy is close to being ruled out. On the other hand, the abundance of $\epsilon=10^{-6}$ is constrained to less than $10^{4}$, which is four to five orders of magnitude than the abundance of Galactic NSs. The weakest constraints in this plot pertain to $\epsilon=10^{-7}$, with a number count of less than $10^{7}$, which is still one to two orders of magnitude smaller than the Galactic NS abundance.
@@
Assuming that the detectable COs are Poisson distributed with a mean $\alpha N_{tot}$, the actual number $n$ that lies within the horizon sphere, with ellipticity $\epsilon$ and frequency $f_{\star}=f_{GW}/2$, has the following discrete probability distribution:
@@
One can equivalently evaluate a spatial-distribution-informed posterior on the upper limit for ellipticity. Assuming that ellipticities are log-uniformly distributed between $[\epsilon_{min},\epsilon_{max}]$, and that the frequencies follow the ATNF catalog, we can evaluate $\alpha$ as a function of $\epsilon_{max}$, for a fixed $\epsilon_{min}$. The posterior on $\epsilon_{max}$ is then given by:
$p(\epsilon_{max}|n=0,N_{tot})\propto\exp(-\alpha(\epsilon_{max})N_{tot})$ (13)
The crucial step in acquiring this posterior is then to construct $\alpha(\epsilon_{max})$. We adopt a Monte-Carlo approach where samples are first drawn from the chosen spatial distribution, the ATNF catalog frequency distribution, and the ellipticities up to an $\epsilon_{max}$. The fraction of samples that lie within the horizon sphere is the desired $\alpha$. Repeating this exercise for a range of $\epsilon_{max}$ enables us to construct $\alpha(\epsilon_{max})$. [^4] The Monte-Carlo approach has its own associated Poisson errors, which have not been accounted for. These errors increase with decreasing ellipticity, where relatively smaller fractions of samples lie within the horizon sphere. Nevertheless, we find that the upper limits quoted on $\epsilon_{max}$ are precise to well within an order of magnitude.
@@ 3 Results
We sample $p(N_{tot}|n=0,\alpha(\epsilon,f_{GW}))$ posteriors and calculate the $90\%$ quantile, $N_{tot}^{90\%}$, across the frequency-ellipticity grid. We plot these upper limits in Figure 2. Each panel in the figure corresponds to one choice of spatial distribution. The contour lines of constant $\log_{10}(N_{tot}^{90\%})$ conservatively suggest a paucity of COs with large ellipticities, potentially (though not strictly) ruling them out. COs with ellipticities $>10^{-4}$ and frequencies $>100$ Hz have $N_{tot}^{90\%}\lesssim 100$, regardless of choice of spatial distribution. Largest allowed ellipticities (Morales & Horowitz, 2022) and frequency $\gtrsim 100$ Hz have $N_{tot}^{90\%}\lesssim 10^{5}$. Indeed, distributions with small scale heights give even more stringent constraints of $N_{tot}^{90\%}\lesssim 10^{4}$, thus strongly limiting the abundance of high-ellipticity Galactic millisecond pulsars.
@@
We plot $p(\epsilon_{max}|n=0,N_{tot}=10^{7,8,9})$ in Fig. 5 for the four choices of spatial distributions. The prior distribution on $\epsilon_{max}$ is taken to be uniform between $10^{-10}$ and $10^{-3}$. Upper limits on $\epsilon_{max}$, at $90\%$ confidence, are also demarcated. Our most conservative estimates suggest that Galactic COs have ellipticities $\lesssim 1.5\times 10^{-7}$, strongly disfavoring eCOs with large ellipticities. This upper limit constraint appears to be slightly stronger than the frequency-dependent upper limits from searches directed toward the Galactic center. However, this improved upper limit is contingent on the choice of frequency and, to a lesser extent, the spatial distribution of COs, and should accordingly be interpreted.
@@ 4 Summary and Outlook
We present a Bayesian framework to infer the abundance of Galactic COs from the non-detection of CGWs, and the frequency-dependent GW amplitude upper limits provided by the searches.  [^5] We refer the reader to other formalisms, (Tenorio, 2023) that discusses constraints on sub-kiloparsec CW sources (e.g. Gravitars), and (Wade et al., 2012) which provides prospective bound on ellipticity and magnetic field parameters with or without detection of CGWs.
@@
Moreover, assuming a log-uniform distribution in ellipticities up to an $\epsilon_{max}$, and an $N_{tot}=10^{7,8,9}$, in addition to the previous assumptions on spatial and frequency distributions, we determine a posterior on $\epsilon_{max}$. The most conservative $90\%$ upper limits suggest $\epsilon_{max}\lesssim 1.5\times 10^{-7}$. This is a stronger (though model-dependent) upper limit than those inferred from searches targeted at the Galactic center (Abbott et al., 2022a).
@@
Figure 5: Posterior on $\epsilon_{max}$, assuming an ATNF-like frequency distribution, four spatial distributions, $N_{tot}=10^{7,8,9}$, and a log-uniform ansatz on the ellipticity distribution with $\epsilon\in[\epsilon=10^{-12},\epsilon_{max}]$. Among all the cases considered, the most conservative upper limit is $\epsilon_{max}\sim 1.5\times 10^{-7}$.
It is interesting to probe the physical parameters of the NS with these constraints. A simple, back-of-the-envelope calculation using Eq. 4 suggests a scarcity of high-magnetic field NSs, with number counts for $B>10^{14}G$ being $N_{tot}^{90\%}\lesssim 10^{6}$. The constraints are even more stringent for high-magnetic field hybrid stars, with $B>10^{14}G$ yielding $N_{tot}^{90\%}\lesssim 10^{2}$ (cf. Eq. 5). However, such statements need to be made more rigorous, since multiple physical mechanisms can impart an ellipticity, and not just magnetic fields. Moreover, spin-down effects from these mechanisms will also need to be accounted for. We leave this for future work.
@@ 5 Acknowledgments
We thank Rodrigo Tenorio and Nathan Johnson-McDaniel for their review and comments on the manuscript. We also thank Debarati Chatterjee, Suprovo Ghosh, Surhud More, Aditya Vijaykumar, Karl Wette, and members of the ICTS Astrophysical Relativity Group, for useful discussions. All computations were performed on the Sarathi Computing Cluster at IUCAA and the Alice Computing Cluster at ICTS-TIFR.

### v3

# Constraining the abundance of spinning deformed Galactic compact objects with continuous gravitational waves

@@ Abstract
Galactic spinning compact objects (COs) with non-zero ellipticity are expected to be sources of continuous gravitational waves (CGWs). Certain classes of hypothetical COs, such as neutron stars with quark cores (hybrid stars), and quark stars, are thought to be capable of sustaining large ellipticities from theoretical considerations. Such exotic COs (eCOs) with large ellipticities and spins should produce CGWs detectable by the current LIGO-Virgo-Kagra GW detector network. Since no detections for CGWs, from searches in LIGO-Virgo data, have so far been reported, we place constraints on the abundance of highly elliptical, rapidly spinning, eCOs in our Galaxy. We formulate a Bayesian framework to place upper limits on the number count $N_{tot}$ of highly deformed Galactic eCOs. We divide our constraints into two classes: an “agnostic” set of upper limits on $N_{tot}$ evaluated on a CGW frequency and ellipticity grid that depend only on the choice of spatial distribution of COs; and a model-dependent set that additionally assumes prior information on the distribution of frequencies. We find that COs with ellipticities $\epsilon\gtrsim 10^{-5}$ have abundance upper limits at $90\%$ confidence, of $N_{tot}^{90\%}\lesssim 100$, and those with $\epsilon\gtrsim 10^{-6}$ have $N_{tot}^{90\%}\lesssim 10^{4}$. We additionally place upper-limits on the ellipticity of Galactic COs informed by our choices of spatial distributions, given different abundances $N_{tot}$.
@@ 1 Introduction
Conversely, the reported non-detection of CGWs from all-sky searches (Abbott et al., 2022b; Steltner et al., 2023; Dergachev & Papa, 2023, 2024) in LV’s O1, O2, O3 data, has strong implications for the abundance of Galactic eCOs. In this letter, we formulate a Bayesian framework to constrain their number count. Specifically, from the CGW (frequency-dependent and sky-averaged) amplitude upper limits (Abbott et al., 2022d), and assuming that the eCOs are Poisson distributed, we construct a Bayesian posterior distribution on the number of COs ($N_{tot}$) from which we place upper limits on the number count at $90\%$ confidence.
@@
We additionally place spatial-distribution-informed upper limits on the ellipticity, assuming CO number counts of $10^{7,8,9}$ (Treves et al., 2000), consistent with the expected abundance of Galactic NSs. Our most conservative estimate yields $\epsilon\lesssim 1.4\times 10^{-7}$.
@@ 2.4 Bayesian framework
Figure 3: Histograms of samples drawn from the posterior on $N_{tot}$ (cf. Eq. 12), marginalized with respect to the ATNF-like frequency distribution. The progenitor and $3$ “Reed” distributions are considered, and the histograms are evaluated for four choices of ellipticities $\epsilon=10^{-7,-6,-5,-4}$. The upper limits for $\epsilon=10^{-5,-4}$ are less than $100$, suggesting that the abundance of such high-ellipticity objects in the Galaxy is close to being ruled out. On the other hand, the abundance of $\epsilon=10^{-6}$ is constrained to less than $10^{4}$, which is four to five orders of magnitude lower than the abundance of Galactic NSs. The weakest constraints in this plot pertain to $\epsilon=10^{-7}$, with a number count of less than $10^{7}$, which is still one to two orders of magnitude smaller than the Galactic NS abundance.
@@
The actual number $n$ that lies within the horizon sphere, with ellipticity $\epsilon$ and frequency $f_{\star}=f_{GW}/2$, follows a Poisson distribution with a mean $\alpha N_{tot}$:  [^4] Strictly, the detectable COs should conform to a binomial distribution. In the large $N_{tot}$ limit, this binomial distribution can be well approximated to be a Poisson distribution with a mean of successful trials, i.e. total detectable events, $\alpha N_{tot}$.
@@
One can equivalently evaluate a spatial-distribution-informed posterior on the upper limit for ellipticity. Assuming that ellipticities are log-uniformly distributed between $[\epsilon_{min},\epsilon_{max}]$, and that the frequencies follow the ATNF catalog, we can construct a posterior distribution on $\epsilon_{max}$ using the following procedure. Similar to Eq(10), the mean fraction of detectable COs given a fixed $N_{tot}$ and a log-uniform distribution on $\epsilon\in[\epsilon_{min},\epsilon_{max}]$ is given by:
[…]
We adopt a Monte-Carlo approach where samples are first drawn from the chosen spatial distribution, the ATNF catalog frequency distribution, and the ellipticities up to an $\epsilon_{max}$. The fraction of samples that lie within the horizon sphere is the desired $\alpha$. Repeating this exercise for a range of $\epsilon_{max}$ enables us to construct $\alpha(\epsilon_{max})$.  [^5] The Monte-Carlo approach has its own associated Poisson errors, which have not been accounted for. These errors increase with decreasing ellipticity, where relatively smaller fractions of samples lie within the horizon sphere.
Thus, analogous to Eq(11), the likelihood function of $\alpha(\epsilon_{max})$ can now be written as:
$p\left(n|N_{tot},\alpha(\epsilon_{max})\right)=\frac{\left(\alpha(\epsilon_{max})N_{tot}\right)^{n}exp\left(-\alpha(\epsilon_{max})N_{tot}\right)}{n!}$ (14)
Re-parameterization of this likelihood function ($\alpha(\epsilon_{max})\rightarrow\epsilon_{max}$) allows us to write the likelihood as  [^6] Note that changing variables in the likelihood function does not require a Jacobian. However, if we had chosen to set a prior on $\alpha(\epsilon_{max})$, we would need to transform that prior into one for $\epsilon_{max}$ using an appropriate Jacobian for the variable transformation. Instead, we opt for a uniform prior directly on $\epsilon_{max}$, rather than setting a prior on $\alpha(\epsilon_{max})$.:
$p(n|N_{tot},\alpha(\epsilon_{max}))=p(n|N_{tot},\epsilon_{max})$ (15)
Now, using Bayes’ theorem, and asserting as before that exactly $n=0$ detections were made, we acquire the posterior on $\epsilon_{max}$ (cf. Eq. 14):
$p(\epsilon_{max}|n=0,N_{tot})$ $\propto p(n=0|N_{tot},\epsilon_{max})p(\epsilon_{max})$
$\propto\exp(-\alpha(\epsilon_{max})N_{tot})$ (16)
@@ 3 Results
We sample $p(N_{tot}|n=0,\alpha(\epsilon,f_{GW}))$ posteriors and calculate the $90\%$ quantile, $N_{tot}^{90\%}$, across the frequency-ellipticity grid. We plot these upper limits in Figure 2. Each panel in the figure corresponds to one choice of spatial distribution. The contour lines of constant $\log_{10}(N_{tot}^{90\%})$ conservatively suggest a paucity of COs with large ellipticities, potentially (though not strictly) ruling them out. COs with ellipticities $>10^{-4}$ and frequencies $>100$ Hz have $N_{tot}^{90\%}\lesssim 100$, regardless of choice of spatial distribution. Largest allowed NS ellipticities (Morales & Horowitz, 2022) and frequency $\gtrsim 100$ Hz have $N_{tot}^{90\%}\lesssim 10^{5}$. Indeed, distributions with small scale heights give even more stringent constraints of $N_{tot}^{90\%}\lesssim 10^{4}$, thus strongly limiting the abundance of high-ellipticity Galactic millisecond pulsars.
@@
We plot $p(\epsilon_{max}|n=0,N_{tot}=10^{7,8,9})$ in Fig. 5 for the four choices of spatial distributions. We choose an uniformative prior on $\epsilon_{max}$ which we set to be uniform between $10^{-10}$ and $10^{-3}$. Upper limits on $\epsilon_{max}$, at $90\%$ confidence, are also demarcated. Our most conservative estimates suggest that Galactic COs have ellipticities $\lesssim 1.4\times 10^{-7}$, strongly disfavoring eCOs with large ellipticities. This upper limit constraint appears to be slightly stronger than the frequency-dependent upper limits from searches directed toward the Galactic center. However, this improved upper limit is contingent on the choice of frequency and, to a lesser extent, the spatial distribution of COs, and should accordingly be interpreted.
Figure 5: Posterior on $\epsilon_{max}$, assuming an ATNF-like frequency distribution, four spatial distributions, $N_{tot}=10^{7,8,9}$, and a log-uniform ansatz on the ellipticity distribution with $\epsilon\in[\epsilon=10^{-12},\epsilon_{max}]$. Among all the cases considered, the most conservative upper limit is $\epsilon_{max}\sim 1.4\times 10^{-7}$.
@@ 4 Summary and Outlook
We present a Bayesian framework to infer the abundance of Galactic COs from the non-detection of CGWs, and the frequency-dependent GW amplitude upper limits provided by the searches. We refer the reader to other formalisms, (Tenorio, 2023) that discusses constraints on sub-kiloparsec CGW sources (e.g. Gravitars), (Wade et al., 2012) which provides prospective bound on ellipticity and magnetic field parameters with or without detection of CGWs, (Knispel & Allen, 2008) which discusses the expected CGW amplitudes from the galactic population of Gravitars and (Pagliaro et al., 2023) which discusses the prospects of detecting standard NS’s in current and future observing scenarios using realistic simulations.
@@
Moreover, assuming a log-uniform distribution in ellipticities up to an $\epsilon_{max}$, and an $N_{tot}=10^{7,8,9}$, in addition to the previous assumptions on spatial and frequency distributions, we determine a posterior on $\epsilon_{max}$. The most conservative $90\%$ upper limits suggest $\epsilon_{max}\lesssim 1.4\times 10^{-7}$. This is a stronger (though model-dependent) upper limit than those inferred from searches targeted at the Galactic center (Abbott et al., 2022a).
@@
It is interesting to probe the physical parameters of the NS with these constraints. A simple, back-of-the-envelope calculation using Eq. 4 suggests a scarcity of high-magnetic field NSs, with number counts for $B>10^{14}G$ being $N_{tot}^{90\%}\lesssim 10^{6}$. The constraints are even more stringent for high-magnetic field hybrid stars, with $B>10^{14}G$ yielding $N_{tot}^{90\%}\lesssim 10^{2}$ (cf. Eq. 5). However, such statements need to be made more rigorous, since multiple physical mechanisms can impart an ellipticity, and not just magnetic fields. Moreover, spin-down effects from these mechanisms will also need to be accounted for. Using results from other searches for CGWs in O3 data (Steltner et al., 2023; Dergachev & Papa, 2023, 2024), in tandem with the LVK’s results, we plan to put constraints on said physical parameters of Galactic COs in future work.
@@ 5 Acknowledgments
We thank Rodrigo Tenorio and Nathan Johnson-McDaniel for their review and comments on the manuscript. We also thank Debarati Chatterjee, Suprovo Ghosh, Surhud More, Aditya Vijaykumar, Karl Wette, and members of the ICTS Astrophysical Relativity Group, for useful discussions. All computations were performed on the Sarathi Computing Cluster at IUCAA and the Alice Computing Cluster at ICTS-TIFR. AKS and RP acknowledge the support of the Department of Atomic Energy, Government of India, under project No. RTI4001.